hours of four batches of electric lamps:•Batch 1: 1600,1610,1650,1680,1700,1720,1800•Batch 2: 1580,1640,1640,1700,1750•Batch 3: 1460,1550,1600,1620,1640,1660,1740,1820•Batch 4: 1510,1520,1530,1570,1600,1680•Perform an analysis of variance of these data and show that a significance test does not reject their homogeneity


1
Expert's answer
2020-04-08T10:34:44-0400

Using Excel we find "\sigma_1^2=4766.667, \sigma_2^2=4220, \sigma_3^2=12169.64, \sigma_4^2=4136.667."

We have the samples so we will use STDEV.S.

The third batch has the largest variance. The fourth batch has the smallest variance.